OpenStudy (anonymous):

The perimeter of a rectangle is 52. The length is 6 greater than the width. Find the length and width of this rectangle.

OpenStudy (anonymous):

x and (x+6), twice = 52 4x = 40 -> x=10 (width) and y = 10+6 = 16 (length

OpenStudy (anonymous):

Thanks